# Visual-RAG: Benchmarking Text-to-Image Retrieval Augmented Generation for Visual Knowledge Intensive Queries

### v1

@@ 4.2 Query Collection
Our objective is to create queries about distinct visual features of organisms, ensuring that only a small fraction of images within each species display the queried feature. Each question must yield a unique answer and generally hold true for every instance of that species, preserving both specificity and accuracy. Typically, composing such specialized queries would require domain expertise and a labour-intensive review of the 200–300 images per species.
Consequently, we employ OpenAI-o1 (OpenAI, 2024) for query generation to replace human experts. As creating meaningful questions requires expert knowledge, we provide the Wikipedia summary and description sections for the model to understand the species. We explicitly instruct the model NOT to generate questions about properties already covered in these text passages, prompting it to focus on more specialized visual attributes unlikely to appear in standard references. The reasoning and self-validating ability of o1 enables it to follow the complex instructions, leading to high-quality and diverse prototype queries. The model is tasked with generating both a set of candidate queries and their corresponding visual features. A refined human annotator filtering and rewriting process is then applied based on these visual features, to ensure query validity [^2] More details about the query generation prompt and human filtering process can be found in Appendix C, D..
@@
The question generation process cannot guarantee compliance with the previously mentioned requirement, i.e., clue images containing the answer should remain a minority in the set; also, query should not be so difficult that no clue image exists in the corpus. We designate a threshold clue image rate, that at most 20% of images in a species level sub-corpus can be clue.
@@
Firstly, we employ an open-sourced MLLM [^3] Llama-3.2-11B-Vision-Instruct was used. Prompt template can be found in Appendix D as a coarse-level filter. For each query, the model iterates over the sub-corpus one image at a time, binary labelling the images by whether the referenced feature is present. These labels are used to roughly filter out queries that exceed the 20% clue-image threshold. Our rationale is that if an unusually large fraction of images exhibit a visual feature, it is likely prominent for the MLLM to detect and generate labels that are reliable enough.
Secondly, we recruit volunteer university students as annotators to label the images. These annotators identify whether each image provides evidence for the question and thereby confirm whether (a) at least one valid clue image exists, excluding overly difficult queries having no clue image and (b) fewer than 20% of the images are clue, excluding remaining “easy” queries not recognized by the coarse-filter. At the same time, the annotators provide the ground-truth answers for each query. This two-stage process ensures our final dataset meets both the clue-image minority requirement and the fundamental criterion that each query should be indeed answerable by evidence clue image within the corpus.
In the current version of the benchmark, we initially select around 1000 questions that passed the coarse filter. After human annotation, 400 of these questions satisfy the clue-image rate threshold and are confirmed answerable with at least one valid clue image.
5 Experimental Setup
5.1 Retrieval Evaluation
For retrieval, we evaluate the performance with Normalized Discounted Cumulative Gain (NDCG) and hit rate – the probability that at least one clue image is found among the top-$k$ retrieved samples (NDCG@k and hit@k).
@@ 5.1 Retrieval Evaluation
Due to the challenge of complex domain-specific queries, it is significantly difficult for the retrievers to search through the entire iNat21 corpus (2.7M images) while the query is about a sophisticated feature for a single species. To enhance retrieval effectiveness and emphasize the more fine-grained and distinguishing features of a species, we narrow the search space by limiting the retrieval corpus to a set of 200–300 images corresponding to the species for each query. In all following experiments, we adopt the commonly used retriever clip-vit-large-patch14-336 (Radford et al., 2021) for image retrieval. The search indexes are constructed with the FAISS library (Douze et al., 2024) using flat inner product index.
5.2 Generation Evaluation
@@
Previous work has relied predominantly on exact-match or recall-based metrics to assess answer correctness. However, such methods are prone to overlooking partial hallucinations blended with ground-truth answers. For example, if the true answer is “black and white” but the model predicts “black and white with yellow dots,” the spurious mention of “yellow dots” remains undetected under exact-match criteria. To address this limitation, we employ GPT-4o as a more nuanced evaluator. Detailed evaluation prompt template can be found in Appendix D. A series of works, including but not limited to Kamalloo et al. (2023); Zheng et al. (2023); Huang et al. (2024), have demonstrated that general-purpose LLMs can reliably judge open-ended QA responses. We also manually validate the automatic evaluation results on a sample of 100 predicted answers, achieving a 94% accuracy rate.
@@
In addition, we report the ROUGE score (Lin, 2004), a metric commonly used in summarization evaluation, as a relaxed version of exact match score. The ROUGE scores are omitted in result analysis section, and can be found in Appendix E.
5.3 Multimodal LLMs
@@
We evaluate 3 advanced proprietary models: GPT-4o (OpenAI, 2025), Gemini-1.5-Pro (Google, 2025), Claude-3.5-Sonnet (Anthropic, 2025); and 5 recent open-sourced models that can process multiple images, of sizes ranging from 4B to 12B: Phi-3.5-vision-instruct (4B, Microsoft, 2024), Qwen2-VL-7B-Instruct (Wang et al., 2024), InternVL-2.5-8B (Chen et al., 2024), Llama-3.2-11B-Vision-Instruct (Meta Llama Team, 2024), and Pixtral-12B (Mistral AI, 2024). We use the default setting given by each model to perform question answering.
6 Results
6.1 Text-to-Image Retrieval
Metric @1 @5 @10 @20 @30
NDCG 26.06 26.46 30.67 38.50 45.80
Hit Rate 26.06 52.73 64.55 72.73 77.87
Hit Count 0.26 1.12 2.08 3.60 4.85
@@ 6.1 Text-to-Image Retrieval
Table 2: Retrieval Results using the sub-corpus of 200-300 images for each query. Even within the small species level corpus, the model struggles with our challenging text-to-image retrieval task.
Successfully retrieving clue images that contain the required visual knowledge is critical to the overall performance of RAG systems in answering the questions. The hit rate highlights the difficulty of text-to-image retrieval in our Visual-RAG benchmark. Even within the small sub-corpus of 200 - 300 images, there is still 22% chance that no clue image can be retrieved within the top-30 samples. The NDCG score also indicates the challenge of this task, i.e., the retrieved clue images are not assigned top similarity scores comparing to the hard negative images.
6.2 Retrieval-Augmented Generation
Model Phi3.5-V Qwen2VL InternVL2.5 Pixtral Llama3.2-V GPT-4o Gemini Claude
@@ 6.2 Retrieval-Augmented Generation
Zero-shot 27.57 32.04 26.17 30.21 35.80 47.21 42.78 42.43
Oracle 32.91 43.19 38.58 37.79 41.51 37.65 18.56 32.19
Non-clue 27.77 33.27 31.53 34.94 31.63 16.61 6.25 8.92
Retrieved Top-K
K=1 33.47 39.24 34.13 37.45 34.59 28.09 19.42 18.53
3 31.14 38.05 32.93 39.44 36.72 35.26 45.48 28.49
5 35.66 40.24 35.52 37.25 39.51 40.84 45.35 30.88
7 33.67 40.44 35.33 37.25 36.65 43.82 46.81 35.86
10 33.47 39.04 37.18 36.85 37.18 44.02 48.40 38.84
15 33.86 39.64 34.39 37.25 35.99 44.42 51.20 45.62
20 32.27 40.84 36.59 - - 47.61 53.19 42.23
Oracle 32.91 43.19 38.58 37.79 41.51 37.65 18.56 32.79
One-in-K
K=3 33.03 40.72 36.72 37.21 37.00 43.90 45.64 32.87
5 33.97 39.68 35.22 39.70 35.92 45.90 46.35 35.98
7 33.35 39.36 34.80 35.66 36.53 45.31 45.46 36.60
10 32.27 39.16 35.61 36.57 34.51 46.14 46.90 37.63
15 32.23 39.40 34.99 37.32 34.59 46.54 46.58 37.86
20 32.39 37.73 34.08 - - 45.62 49.57 35.43
Oracle 32.91 43.19 38.58 37.79 41.51 37.65 18.56 32.79
Table 3: Main experiment results. The coloured cells shows the difference with oracle score, best performing $k$ values for each model is highlighted with boldface.
@@
To assess whether image can be leveraged as evidence in the RAG system, we conduct baseline experiments and compare performance under three conditions: 1) Zero-shot: directly prompting MLLMs with questions without retrieval; 2) Oracle: augmenting the question with a gold clue image; 3) Non-clue: augmenting each question with a non-clue image within the same species, which does not contain visual knowledge relevant to the question [^4] For 2) and 3), the results reported are the average of 5 runs with different clue / non-clue images selected as augmentation., as shown in the first section of Table 3.
All five open-sourced models exhibit performance gains of up to 15 points when using oracle image as an augmentation. This indicates the necessity of including relevant visual knowledge to answer the questions. We conclude that for open-sourced models, a single relevant image can serve as good evidence to enhance RAG performance.
Whereas interestingly, oracle performance for proprietary models drops below the zero-shot setting. We hypothesize that the single-image clue provides insufficient evidence to establish the correct visual relationship, causing the model to respond conservatively with “I cannot decide”. These conservative responses count as failing to produce an answer, thus reducing the overall scores. Specifically, such conservative response occurs in 15.8% of samples for GPT-4o, 7.4% for Gemini, and 16.4% for Claude. Meanwhile, results for the Non-clue setting indicate that the models effectively identify these non-clue images as not containing visual knowledge. Specifically, all proprietary models frequently respond with “I don’t know” with non-clue images, leading to a substantial drop in scores comparing to the zero-shot and oracle setting.
Can MLLMs distinguish the clue from retrieved hard negatives?
@@
To explore this realistic RAG scenario, we provide the top-$k$ retrieved images, derived from Section 6.1, as augmentation to the MLLMs [^5] For pixtral and Llama3.2-V, $k=20$ was not evaluated due to hardware limitation.. The retrieved images are sorted in descending order based on their similarity scores. As shown in the second section of Table 3, the open-source models generally lag behind their single-oracle scores, indicating limited multi-image processing capabilities. Most open-source models struggle to isolate the clue image and effectively leverage its visual information when it is mingled with hard negatives.
In contrast, all three proprietary models exhibit a generally monotonic increase in performance as $k$ increases, except Claude which decreases at $k=20$. Notably, the performance at $k=5$ begins to surpass oracle baselines, continuing to improve along with higher $k$ values. We attribute this improvement to the comparative evaluation enabled by providing both clue and non-clue images. Compared to a single clue image, the top-$k$ RAG setting allows the models to contrast relevant and irrelevant visual features, leading to greater confidence in extracting knowledge from the clue images, whereas a single clue image may be insufficient to fully establish its relationship to the question. However, the proprietary models fail to exceed their zero-shot scores until fairly large $k$ values, suggesting that interpreting complex visual cues remains challenging, and they may even struggle with reconciling wrongly interpreted visual knowledge with their prior knowledge gained through pre-training.
As indicated by the retrieval results, even at $k=20$, there is 27% chance that no clue image is included in the retrieved set. This likely explains the low QA performance at smaller $k$ values, where failing to retrieve the clue image entirely negates any benefit of visual augmentation. To investigate further, we conduct an additional experiment, constructing image sets in which exactly one clue image and $k-1$ non-clue images are provided [^6] Both clue and non-clue images were randomly selected across 5 different combinations, results reported are the average of 5 runs for each $k$.. The clue image is positioned at the beginning of the $k$ images to simulate the perfect retrieval scenario.
Demonstrated in the last section of Table 3, the open-sourced models now exhibit decreasing performance as $k$ increases, further strengthen the conclusion of their limited capacity for handling multiple images simultaneously. By contrast, all three proprietary models surpass their single-image oracle baseline from $k=3$ onward and maintain near-constant performance at higher $k$ values. This suggests that proprietary models are capable of isolating the clue image from irrelevant images, regardless of the number of input images [^7] Additional experiments on randomly positioning the clue image within the set is also carried out. The proprietary models keeps the flat performance curve against $k$, indicating there is no positional bias..
Model Zeroshot Score Augmenting Text
Phi3.5-V 27.57 26.02
Qwen2VL 32.04 25.61
InternVL2.5 26.17 19.70
Pixtral 30.21 19.74
Llama3.2-V 35.80 29.73
Table 4: Results when augmenting retrieved Wikipedia text as evidence.
6.3 Ablation
Can the answers for Visual-RAG be found in Wikipedia?
We construct a small textual knowledge base from Wikipedia articles for the 10k species in iNat21 dataset. This minimum knowledge base comprising approximately 97,000 chunks, each consisting of $\sim$150 tokens. The detailed collection process is described in Appendix B. For text retrieval, we use a pre-trained dense text retriever, DRAGON Plus (Lin et al., 2023). Given that the ground-truth labels for relevant text chunks are not available, we only assess whether the retrieved text chunks correspond to the correct Wikipedia page, and the retriever achieves HitRate@10 of 97%. Next, we present the QA performance across open-sourced models augmented with top-10 retrieved textual knowledge chunks. As shown in Table 4, augmenting the question with textual knowledge from Wikipedia does not improve answer quality for most models. Instead, in numerous cases, models fail to locate relevant information within the augmented text and output “I don’t know”. This behaviour in turn leads to scores lower than the zero-shot setting.
These experiments provide strong evidence that the knowledge required to answer the questions related to visual features in our benchmark is unlikely to be contained in textual knowledge bases.
Is More Clues Always Better?
We observed that for proprietary models, top-$k$ scores at higher $k$ values excel their 1-in-$k$ score. Similarly, the open-source models also achieve their best performance at larger $k$ value in top-$k$ setting than in 1-in-$k$. A likely explanation is that the expected number of clue images retrieved in top-$k$ is substantially higher than when only a single clue image is guaranteed. We in turn ablate with effect of multiple clue images. Fixing $k=10$, we experiment with varying numbers of clue images with $n\in\{1,2,3,4,5\}$ ($n$-in-10). We select queries that contain at least 5 clues to enable this experiment. As shown in Figure 3, all models exhibit increasing performance as $n$ grows, aligning with the intuition that additional clue images provide more robust visual evidence.
Figure 3: Impact from number of clue images
Hard-negatives can be helpful; how about easy-negatives?
As an ablation to the previous experiment on whether multiple images benefit the overall QA performance, we now mix one clue image with $k-1$ highly irrelevant images from subsets that are not the query species. The results indicate that both open-source and proprietary models benefit more from hard negatives coming from the same species compared with easy negatives coming from different species. This further strengthen our conclusion that hard negatives are helpful in contrasting the queried features to inform the correct answer. Detailed results can be found in Table 9 in Appendix.

### v2

@@ 4.2 Query Collection
Our goal is to create text queries about distinct visual features of organisms such that, for a given species, only a small fraction of images exhibit the queried feature. Each query must admit a unique answer or answer set that generally holds at the species level (category‑level fact), rather than instance‑specific properties.
Constructing such queries by hand would require domain expertise and careful inspection of roughly 200–300 images per species. To scale this process, we employ OpenAI‑o1 as a query generator. For each candidate species, we provide its Wikipedia Summary and Description sections to the model to convey basic context. We explicitly instruct the model NOT to generate questions about properties already described in these passages, pushing it toward specialized visual attributes that are unlikely to appear in standard references. The model outputs (i) a set of candidate questions and (ii) their associated visual feature descriptors (e.g., “hindwing color”, “presence of mottled pattern”, etc.).
To reduce the chance that answers can be trivially recovered from text corpora, we bias species selection toward those that are relatively under‑represented in Wikipedia. A subsequent refined human filtering and rewriting process (Appendix C, D) validates the linguistic form of each query and the feature descriptors.
@@
Our curation process enforces two constraints: (C1) clue images (images that clearly display the queried feature with the required view/pose) must be a minority within the species‑specific sub‑corpus; and (C2) each query must be answerable by at least one clue image. Formally, for a query $q$ over species $s(q)$ with image set $\mathcal{D}_{s(q)}$, let $\mathcal{I}_{c}\subset\mathcal{D}_{s(q)}$ denote the clue set and define the clue rate $\rho_{q}\triangleq\frac{|\mathcal{I}_{c}|}{|\mathcal{D}_{s(q)}|}$. We retain only queries within threshold $0<\rho_{q}\leq 0.25$.
Stage 1: MLLM‑based coarse filter.
@@
We use an open‑source MLLM [^2] Llama-3.2-11B-Vision-Instruct. Prompt template can be found in Appendix D. as an image‑level presence / absence classifier for the referenced feature. For each query $q$, the model iterates over $\mathcal{D}_{s(q)}$ one image at a time and predicts the binary label indicating whether the feature is visually present. From these labels we obtain an estimate clue rate $\hat{\rho}_{q}$ and discard queries with $\hat{\rho}_{q}>0.25$. The premise is that when a queried feature is unusually common, it is sufficiently prominent for the MLLM to detect reliably, making the coarse estimate adequate for pre‑screening.
Stage 2: Human annotation and answer adjudication.
We recruit volunteer university students as annotators to (a) label images by whether the queried feature is visibly present and (b) provide the ground‑truth answer for the query $q$. This step confirms that (i) at least one valid clue image exists and (ii) $\rho_{q}\leq 0.25$, thereby removing queries that slipped past the coarse filter. We take the human labels as ground truth for $\mathcal{I}_{c}$ and compute the final $\rho_{q}$ accordingly. Approximately 1,0001{,}000 questions pass the coarse MLLM filter. After human verification, 374 questions satisfy the clue‑rate criterion ($0<\rho_{q}\leq 0.25$) and are confirmed answerable by at least one clue image.
5 Experiments
5.1 Experimental Setups
Retrieval Evaluation.
For retrieval, we evaluate the performance with Recall score, Normalized Discounted Cumulative Gain (NDCG), Hit rate and Hit Count. Recall@$k$ measures the fraction of all ground-truth clue images that appear within the top-$k$ results, reflecting overall coverage. NDCG@$k$ rewards correct images that appear higher in the ranking by discounting relevance logarithmically with position, thus capturing ranking quality. Hit@$k$ is a binary success indicator for probability of at least one clue is retrieved in the top-$k$ images, while Hit Count@$k$ tallies how many clue images are found in that cut-off, offering an intuitive sense of how much useful evidence the retriever supplies.
@@ 5.1 Experimental Setups
Due to the challenge of complex domain-specific queries, it is significantly difficult for the retrievers to search through the entire iNat21 corpus (2.7M images) while the query is about a sophisticated feature for a single species. To enhance retrieval effectiveness and emphasize the more fine-grained and distinguishing features of a species, we narrow the search space by limiting the retrieval corpus to $\mathcal{D}_{s(q)}$, the set of 200–300 images corresponding to the species for each query.
Dense Retrievers.
Three dense retriever models are evaluated: (1) The commonly used multimodal retriever CLIP (clip-vit-large-patch14-336, Radford et al. 2021). (2) A recent CLIP variant, BGE (BGE-VL-large, Zhou et al. 2024. (3) MLLM-based retriever, VLM2Vec-qwen2VL-2B (Jiang et al. 2025). The search indexes are constructed with the FAISS library (Douze et al. 2024) using flat inner product index. RAG experiments in following sections adopt images retrieved by CLIP, as CLIP is being adopted as the de-facto retriever in previous multimodal RAG works.
Generation Evaluation.
@@
Previous work has relied predominantly on exact-match or recall-based metrics to assess answer correctness. However, such methods are prone to overlooking partial hallucinations blended with ground-truth answers. For example, if the true answer is “black and white” but the model predicts “black and white with yellow dots”, the spurious mention of “yellow dots” remains undetected under exact-match criteria.
To address this limitation, we adopt LLM-as-Judge, employing OpenAI o4-mini as a more nuanced evaluator. We retain all annotator-provided answers as valid ground-truth candidates, and specify in instruction prompt to consider alternative expressions of features. Detailed evaluation prompt template can be found in Appendix D. A series of works, including but not limited to Kamalloo et al. (2023); Zheng et al. (2023); Huang et al. (2024), have demonstrated that general-purpose LLMs can reliably judge open-ended QA responses. We also manually evaluated 200 sampled predicted answers, and the LLM labels yielded alignment score of $F_{1}=0.92$ with human annotator. The confusion matrix can be found in Appendix  E.
@@
In addition, we report the ROUGE score (Lin 2004), a metric commonly used in summarization evaluation, as a relaxed version of exact match score. The ROUGE scores are omitted in result analysis section, and can be found in Appendix E.
Multimodal LLMs.
@@
We evaluate three flagship proprietary models: GPT-4o (OpenAI 2025), Gemini-2.5-Pro (Google 2025), Claude-Sonnet-4 (Anthropic 2025); [^3] GPT-4o does not internally implement long Chain-of-Though reasoning, while Gemini-2.5-Pro and Claude-Sonnet-4 implement long CoT reasoning (named Thinking Mode) by default. For fair comparison, we disabled Thinking Mode in Claude; Gemini-2.5-Pro does not support full disabling Thinking Mode, we set the thinking token budget to the lowest possible value 128. and five recent open-source models that can process multiple images, of sizes ranging from 5B to 12B: Phi-4-Multimodal-Instruct (5.6B, Microsoft 2025), Qwen2.5-VL-7B-Instruct (Bai et al. 2025), InternVL-3-8B (Zhu et al. 2025), Llama-3.2-11B-Vision-Instruct (Meta Llama Team 2024), and Pixtral-12B (Mistral AI 2024). We use the default inference parameters given by each model in generation.
CLIP @1 @5 @10 @20 @30
Recall 2.81 10.26 16.70 25.54 33.16
NDCG 24.33 25.97 30.98 39.21 46.25
Hit 24.33 53.74 67.65 77.54 82.62
Hit Count 0.24 1.01 1.84 3.18 4.38
BGE @1 @5 @10 @20 @30
Recall 3.63 12.43 18.81 28.64 35.97
NDCG 26.74 29.15 33.4 41.83 48.84
Hit 26.74 58.56 68.98 79.14 83.96
Hit Count 0.27 1.17 2.04 3.43 4.60
VLM2Vec @1 @5 @10 @20 @30
Recall 2.91 10.01 16.03 26.37 33.84
NDCG 24.87 26.33 30.09 39.13 45.93
Hit 24.87 51.87 62.83 78.34 82.35
Hit Count 0.25 1.10 1.91 3.33 4.49
@@
Table 2: Retrieval results using the sub-corpus of 200-300 images for each query. Even within the small, species level corpus, the model struggles with our challenging text-to-image retrieval task.
5.2 Results of Text-to-Image Retrieval
Model Phi4-$M$ Qwen2.5VL InternVL3 Pixtral Llama3.2-V GPT-4o Gemini Claude
@@ 5.2 Results of Text-to-Image Retrieval
Zero-shot (no image) 35.16 38.90 39.17 41.71 32.35 53.74 60.43 54.28
GT clue (1 image) 45.04 41.79 43.69 47.11 47.81 59.81 62.88 56.79
Non-clue (1 image) 34.76 30.08 29.28 42.74 40.37 14.97 17.11 21.39
Top-K Retrieval-Augmented Generation
k=1 39.17 41.98 37.57 42.11 44.25 24.06 32.22 35.03
3 39.30 44.39 41.18 39.71 46.79 41.18 48.53 45.32
5 37.70 46.93 39.97 41.71 43.80 47.86 54.95 50.40
7 41.44 45.99 42.65 41.44 43.98 51.34 55.35 52.14
10 41.44 48.26 42.11 42.65 43.42 49.73 57.62 53.88
15 41.04 49.73 41.04 44.39 44.92 50.80 60.03 55.48
20 41.31 50.53 41.31 - - 52.67 61.50 57.35
One-in-K Augmented Generation
k=3 41.92 46.85 41.06 47.11 46.73 48.95 59.57 48.05
5 40.13 48.04 38.44 46.68 45.19 53.22 60.70 49.86
7 40.80 47.78 39.76 46.68 44.20 55.36 61.53 51.01
10 39.08 47.43 38.50 44.83 41.90 56.25 62.70 51.69
15 40.40 49.57 39.14 43.95 41.12 56.36 64.01 52.67
20 40.83 48.55 39.09 - - 56.50 63.47 52.14
Table 3: Main experiment results. The coloured cells shows the difference with zero-shot score, pink cells indicate performance under zero-shot baseline, light green cells indicate performance over zero-shot, but lower than GT, and green cells indicate outperforming GT. All models benefit from ground-truth clue image as augmentation.
Overall difficulty.
Successfully retrieving clue images that contain the required visual knowledge is critical to the overall performance of RAG systems in answering the questions. As shown in Table 2, across the three retrievers, performance is tightly clustered, with BGE slightly ahead. Even when restricting the search space to 200–300 images per species, retrieval remains challenging: across all retrievers, there is $\sim$ 17-18% chance that no clue image appears within the top‑30. These results indicate that (i) clue images are genuinely rare and (ii) clue images, if retrieved, are not ranked top by similarity scores. The sub-optimal performance of retrievers poses a further challenge for MLLMs, requiring models to tackle with the task under noisy image context.
Takeaway.
Substantial performance gap remains to be closed for cross-modal retriever. Performance remains far from saturated even at $k=30$, downstream gains will depend not only on the MLLM’s visual reasoning but also on the retriever exposing more clues, and earlier in the ranking.
5.3 Results of Visual Retrieval-Augmented Generation
@@
To assess whether image can be leveraged as evidence in the RAG system, we conduct baseline experiments and compare performance under three conditions: 1) Zero-shot: directly prompting MLLMs with text questions without image retrieval; 2) GT clue: augmenting the question with one ground-truth clue image; [^4] Results reported are the average of 5 runs with different random clue image selected when possible. 3) Non-clue: augmenting each question with a non-clue image within the same species, which does not contain visual knowledge relevant to the question, as shown in the first section of Table 3.
GT clue setting establishes an upper bound on how images can contribute in RAG. Relative to zero-shot baseline, supplying a single ground-truth clue image raises performance for all models tested. Open-source models receive up to 15 accuracy scores gain, confirming that augmenting relevant visual evidence can substantially compensate for gaps in their pre-trained prior knowledge. The three proprietary models, while having strong zero-shot baseline performance, also improved with augmenting clue image for 2-6 points in accuracy. Non-clue image, conversely, degrades performance of most models, underscoring that the benefit comes from evidence encoded in the clue itself, not from the mere presence of an image.
Takeaway: A single, representative clue image reliably improves performance, confirming visual evidence in our benchmark is informative and influential for RAG.
Does realistic visual RAG benefit MLLMs?
@@
To explore this realistic RAG scenario, we provide the top-$k$ retrieved images, [^5] For Pixtral and Llama3.2-V, $k=20$ was not evaluated due to hardware limitation. derived from Section 5.2, as augmentation to the MLLMs. The retrieved images are sorted in descending order based on their similarity scores.
Intuitively, more retrieved images would lead to better performance, as probability of retrieving clue image increase with $k$. As shown in the second section of Table 3, while Qwen and the three proprietary models follow the expected trend, other open-source models’ performance saturates and plateaus at medium $k$ (5 - 10), including more retrieved images offers little benefit and sometimes slightly hinders the performance.
Two intertwined factors plausibly explain the divergence: the open-source models have limited multiple-image processing capacity, as $k$ increase beyond 15, they struggle with effectively extracting related visual information; in the meantime, while larger $k$ brings higher probability of retrieving clue images, the number of negative non-clue images occupy larger portion of the image set. This further dilutes the clue signal and magnifies the impact of their restricted multi-image understanding ability.
Takeaway: Under realistic RAG setting, most of the evaluated models fall noticeably short of the GT clue upper bound, revealing current MLLMs’ limitation in efficiently utilizing retrieved image context. However, Qwen stands as a powerful exception, possessing robust multimodal RAG ability.
Does a guaranteed clue outweigh the noise of additional non-clue images?
We notice that top-$k$ RAG underperform the GT clue setting for most models. Such results conflate two error sources: (i) the retriever may fail to return any clue image, or (ii) clue present, but models fail to extract information alongside visually similar non-clue distractors. To investigate further, we conduct an additional experiment named 1-in-$k$, constructing image sets in which exactly one clue image and k1k-1 non-clue images are provided,6665 runs for different random images combinations at each $k$. isolating the impact from imperfect retrieval. The clue image is positioned at the beginning of the $k$ images, emulating perfect retrieval scenario.
Demonstrated in the last section of Table 3, most open-source models except Qwen now lose accuracy as $k$ increases. Proprietary models, in stark contrast, keep improving monotonically with $k$, while GPT-4o and Gemini surpass their top-$k$ RAG performance. Hence, the plateau observed in earlier RAG experiments for open-source models is not merely a retrieval problem: even when the clue is guaranteed, models are still prone to more distractors. Conversely, the unsatisfactory performance of proprietary models at lower $k$ under top-$k$ RAG can be confirmed to be attributable from missing clue image retrieved.
Takeaway: With a guaranteed clue image, proprietary models remain robust to additional non-clue distractors, underscoring their ability to extract visual evidence; whereas open-source models are markedly distractor-sensitive.
Figure 3: gCUE metrics by models and $k$.
How efficiently do MLLMs utilize clue and how robust are they against distractors?
With the 1-in-$k$ results, we propose a quantitatively metric named generalized clue utilization efficiency (gCUE). Specifically, we integrate zero-shot and non-clue baselines via $B^{(\lambda)}=\lambda A_{Z}+(1-\lambda)A_{NC}$, where $\lambda\in[0,1]$, $A_{Z}$ denotes zero-shot accuracy, $A_{NC}$ denotes non-clue accuracy, $A_{GT}$ denotes GT clue accuracy and $A_{k}$ denotes 1-in-$k$ accuracy at $k$. The non-clue baseline $A_{NC}$ is introduced as an approximation of the impact on performance from non-clue images. We define:
$\mathrm{gCUE}(k;\lambda)=\frac{A_{k}-B^{(\lambda)}}{A_{GT}-B^{(\lambda)}}.$
$\mathrm{gCUE}$ therefore quantifies the remaining benefit of the clue image over baseline among k1k-1 distractors, compared with sole clue image scenario. It describes how effectively the model utilizes the visual clue and how it is prone to distractors. $\mathrm{gCUE}=1$ implies model attaining full GT clue benefit, $\mathrm{gCUE}<1$ indicates under-utilization of the clue and vulnerability to distractors, and $\mathrm{gCUE}>1$ suggests model gaining synergy with non-clues. By default we set $\lambda=0.5$.
As shown in Figure 3, Qwen demonstrates strong synergy and distractor tolerance across $k$. Other open-source models generally display a trend of decline. Their ability to exploit the single clue erodes, suggesting limited capacity for multi-image aggregation and greater vulnerability to a rising distractor load. Whereas proprietary models start with low $\mathrm{gCUE}$ at small $k$, implying that a handful of non-clue images can initially overwhelm the clue despite being positioned at front. Yet they surprisingly recover as more non-clue context being provided. One plausible explanation is that the proprietary models and Qwen implicitly contrast relevant and irrelevant visual features among the noisy image set. As more non-clues are added, the distinctive features of the clue become easier to isolate, eventually lead to confidence in extracting knowledge from the clue image. A detailed causal analysis of this effect is beyond the scope of the present paper and we leave it to future work.
Takeaway: $\mathrm{gCUE}$ metric provides an interpretable metric apart from raw accuracy. Open-source models except Qwen continue to exhibit inverted dynamics comparing to proprietary models on clue utilization efficiency.